# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

### v2

@@ Introduction
Let $k\leq n^{\prime}/2$. Consider using the GSEMO to optimize the $m$OJZJ problem. Then after an expected number of at most $eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ iterations (or $1+eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ fitness evaluations), the full Pareto front is covered.
Reduced Impact of Stochastic Population Update
Bian, Zhou, Li, and Qian (2023) proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains. More precisely, it was proven that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the bi-objective OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice. Hence, for $k=\omega(\log n)$, a super-polynomial speed-up was shown.
In this section, we extend the analysis of (Bian et al. 2023) to $m$ objectives. Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update. As in (Bian et al. 2023), we have no proven tight lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
Stochastic Population Update
The rough idea of the stochastic selection proposed by Bian et al. (2023) is that a random half of the individuals survive into the next generation regardless of their quality. The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA. This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently (Lehre and Özcan 2013; Lissovoi, Oliveto, and Warwicker 2023; Doerr et al. 2023).
More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement. The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only. We note that with this mechanism, even the worst solution enters the next generation with probability at least $1/2$.
Runtime
@@
Consider using the SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem. For any iteration $t$ and $x\in R_{t}$, we have that there is a $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
@@
Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. If at some time $t$ the combined parent and offspring population $R_{t}$ of the SMS-EMOA contains some solution $x$, then at any later time $s>t$ the parent population $P_{s}$ contains a solution $y$ such that $y\succeq x$.
Since the proofs of Lemmas 5 and 6 for the classic SMS-EMOA mostly relied on elementary properties of standard bit-wise mutation and on the survival guarantee of Lemma 4, we can now use the survival guarantee of Lemma 11 to obtain analogous results for the SMS-EMOA with stochastic selection (at the price of requiring essentially twice the population size). This yields the following estimates for the time to obtain at least one inner Pareto optimum and the time to cover all inner Pareto front points starting from a population with at least one inner Pareto optimum.
@@
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
  • after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
  • after another at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered. Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks they contain. The difference here is that we consider the effect of the gap points, similar to Bian et al. (2023).
@@
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Combining Lemmas 12 and 13, we have the runtime of the SMS-EMOA with stochastic population update in the following theorem.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq 2(M+1)$ and stochastic population update to optimize the $m$OJZJ problem. Then after at most $e\mu(mk)^{k}(1+\ln m)+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Comparing the runtime guarantees or Theorem 8 (classic SMS-EMOA) and Theorem 14 (SMS-EMOA with stochastic selection), we see that stochastic selection can at most lead to a speed-up by a factor of order $2^{k/2}/\mu k^{1/2}$. Now $\mu$ is at least $\Omega(M)=\Omega((2n/m-2k+3)^{m/2})$. Consequently, the advantage of stochastic selection is rapidly decreasing with growing numbers of objectives and vanishes, e.g., when $m\geq k$.
Heavy-Tailed Mutation Helps
In the previous section, we saw that the advantages of a stochastic population update do not well generalize from bi-objective optimization to many-objective optimization. We now regard another design choice that so far was only analyzed in bi-objective optimization, namely a heavy-tailed mutation operator. We shall prove that the $k^{\Omega(k)}$ factor speed-up observed in bi-objective optimization extends to many objectives.
Heavy-Tailed Mutation
Different from the standard bit-wise mutation operator, which flips each bit independently with probability $1/n$, the heavy-tailed mutation operator proposed in (Doerr et al. 2017) flips each bit independently with probability $\alpha/n$, where $\alpha$ follows a power-law distribution with parameter $\beta$. The number $\alpha$ is sampled anew in each application of the heavy-tailed mutation operator. The underlying power-law is defined as follows.
@@
Let $n\in\mathbb{N}$ and $\beta>1$. We say $\alpha$ follows a power-law distribution with (negative) exponent $\beta$ if for all $i\in[1..n/2]$, we have $\Pr[\alpha=i]=\left(\sum_{j=1}^{n/2}k^{-\beta}\right)^{-1}i^{-\beta}$.
The power-law feature extends to the heavy-tailed mutation operator in the sense that it generates an offspring with Hamming distance $j$ from the parent with probability $\Omega(j^{-\beta})$. This facilitates larger jumps in the search space, and thus escaping from a local optimum.
The heavy-tailed mutation operator has resulted in asymptotic performance gains by a factor of $k^{\Omega(k)}$ for the ${(1+1)}$ EA optimizing single-objective (classic and generalized) Jump functions with gap size $k$ (Doerr et al. 2017; Bambury, Bultel, and Doerr 2021). For the bi-objective OJZJ benchmark, again a speed-up of $k^{\Omega(k)}$ was proven when optimized via the GSEMO (Zheng and Doerr 2023d) and the NSGA-II (Doerr and Qu 2023a). Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choices (Friedrich, Quinzan, and Wagner 2018; Wu, Qian, and Tang 2018; Quinzan et al. 2021; Corus, Oliveto, and Yazdani 2021; Dang et al. 2022; Antipov, Buzdalov, and Doerr 2022; Doerr, Ghannane, and Ibn Brahim 2022; Doerr and Rajabi 2023).
@@
With proof ideas similar to those in Lemma 5 to 7, but noticing the different probabilities of generating solutions with a specific Hamming distance, we obtain the following runtime guarantee.
@@
Let $k\leq n^{\prime}/2$ and $\beta>1$. Consider using SMS-EMOA with $\mu\geq M$ and with heavy-tailed mutation to optimize the $m$OJZJ problem. Then after at most $O(M\mu k^{\beta-0.5-k}n^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
@@
Compared to the runtime guarantee of Theorem 8 for the original SMS-EMOA, the guarantee of Theorem 17 above for the SMS-EMOA with heavy-tailed mutation is by a factor of asymptotically $k^{k+0.5-\beta}$ stronger.
The SMS-EMOA Also Performs Well for Bi-objective Optimization

### v3

@@ 4.3 Runtime of the GSEMO on $m$OJZJ
Let $k\leq n^{\prime}/2$. Consider using the GSEMO to optimize the $m$OJZJ problem. Then after an expected number of at most $e\overline{M}(mk/2)^{k}(1+\ln m)+eM\overline{M}n^{k}$ iterations (or $1+e\overline{M}(mk/2)^{k}(1+\ln m)+eM\overline{M}n^{k}$ fitness evaluations), the full Pareto front is covered.
5 Reduced Impact of Stochastic Population Update
Bian et al. [BZLQ23] proposed a stochastic population update mechanism for the SMS-EMOA and proved, somewhat unexpectedly given the state of the art, that it can lead to significant performance gains. More precisely, it was proven that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the bi-objective OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice. Hence, for $k=\omega(\log n)$, a super-polynomial speed-up was shown.
In this section, we extend the analysis of [BZLQ23] to $m$ objectives. Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update. As in [BZLQ23], we have no proven tight lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
As a side result, we simply and moderately improve the analysis of how the stochastic population update helps to traverse the valley of low-quality solutions of the OJZJ benchmarks. Whereas in [BZLQ23] a complex drift argument was used to show a speed-up of $\mu/2^{k/4}$, we give a simple waiting time argument that yields a speed-up of $\mu/(2^{k}-1)$, both in the bi-objective case of [BZLQ23] and in the general case.
5.1 Stochastic Population Update
The rough idea of the stochastic selection proposed by Bian et al. [BZLQ23] is that a random half of the individuals survive into the next generation regardless of their quality. The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA. This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently [LÖ13, LOW23, DDLS23].
More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement. The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only. We note that with this mechanism, any solution enters the next generation with probability at least $1/2$.
@@ 5.1 Stochastic Population Update
Consider using the SMS-EMOA with stochastic population update to optimize any specific problem. For any iteration $t$ and $x\in R_{t}$, we have that there is a $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
Proof.
We simply calculate the probability of $x$ being removed after the population update. Since $x$ can be removed only if it is selected in $R^{\prime}$, we know that it will be removed with probability at most $\frac{\lfloor(\mu+1)/2\rfloor}{\mu+1}\leq 1/2$, and hence $f(x)$ will be maintained in $P_{t+1}$ with probability at least $1/2$. ∎
@@
Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2\overline{M}+1$ to optimize any specific problem, where $\overline{M}$ is the maximum size of a set of incomparable solutions for this problem. If at some time $t$ the combined parent and offspring population $R_{t}$ of the SMS-EMOA contains some solution $x$, then at any later time $s>t$ the parent population $P_{s}$ contains a solution $y$ such that $y\succeq x$.
Proof.
If $x\notin R^{\prime}$, then it survives automatically into the next generation. If $x\in R^{\prime}$, then it is subject to a selection as in the standard SMS-EMOA. Since $R^{\prime}=\lfloor(\mu+1)/2\rfloor>\overline{M}$, Lemma 4 gives the desired $y$. ∎
5.2 Runtime Analysis for the SMS-EMOA With Stochastic Population Update
We now analyze the runtime of the SMS-EMOA with stochastic population update on the many-objective OJZJ problem. We start by distilling why the SMS-EMOA profits from the stochastic population update and tightening the analysis of [BZLQ23]. Our analysis is also significantly simpler than the previous one, which might easy future works in this direction.
The core of the analysis in [BZLQ23] is proving that the SMS-EMOA can profit from the stochastic population update by accepting solutions in the valley of low-quality solutions of the bi-objective OJZJ benchmark. When extracted from the long proof in [BZLQ23], this would give the following lemma.
Lemma 13 ([BZLQ23]).
Consider using the SMS-EMOA with stochastic population update to optimize the bi-objective OJZJ benchmark. Assume that $\mu\geq 2(M+1)$. Assume that at some time $t$, the population contains an individual with $k$ ones. Then the expected number of additional iterations to generate the all-zero solution is at most $\min\{e\mu n^{k},(e\mu n^{k/2}+1)2e\mu^{2}n^{k/2}/2^{k/4}\}=e\mu n^{k}\min\{1,(1+o(1))2e\mu/2^{k/4}\}$.
The proof of this lemma relies on a technical application of the additive drift theorem of He and Yao [HY01], in particular, with a non-intuitive potential function.
Since the valleys of low-quality solutions of the OJZJ benchmark have the same shape regardless of the number of objectives, we could use the elementary lemmas from the previous subsection to extend the lemma above to the many-objective setting. Since we are not fully satisfied with the complex proof of Lemma 13, we propose an alternative one that appears both simpler and gives a tighter result. We formulate it in the many-objective setting, but note that it includes the bi-objective setting. In terms of the bound, our result replaces the factor of $(1+o(1))2^{-k/4}$, which is the term responsible for the runtime improvement, by a factor of $2^{-k+1}$, so essentially the fourth power of the previous factor.
@@ 5.2 Runtime Analysis for the SMS-EMOA With Stochastic Population Update
Consider using the SMS-EMOA with stochastic population update to optimize the $m$-objective OJZJ benchmark, $m\geq 2$. Assume that $\mu\geq 2\overline{M}+1$. Assume that at some time $t$, the population contains an individual $x$ with $k$ ones in the $i$-th block. Then the expected number of additional iterations to generate an individual $z$ with only zeros in the $i$-th block and with $f_{j}(z)=f_{j}(x)$ for all $j\in[1..m]\setminus\{2i-1,2i\}$ is at most $e\mu n^{k}\min\{1,4e\mu/2^{k}\}$.
Proof.
The proof of this lemma consists of a simple waiting time argument, that is, we compute the probability $p$ that the desired solution $z$ is generated from $x$ in at most two iterations and then bound the expected waiting time by $2/p$.
Let $x$ be as in the assumptions of this lemma. Let $X$ be the set of indices of ones it has in the $i$-th block. Let $Y\subseteq X$. Let $E_{Y}$ be the event that in the first iteration, (i) $x$ is selected as parent for mutation, (ii) it is mutated into an individual $y$ that agrees with $x$ in all bits except for exactly the ones in $Y$, and (iii) this $y$ is taken into the next parent population (despite being dominated by $x$ when $Y\notin\{\emptyset,X\}$), and in the next iteration, (iv) this $y$ is selected as parent for mutation and (v) mutated into a solution $z$ that agrees with $x$ in all bits except for exactly the ones in $X$. The probabilities for these steps to happen as described are $1/\mu$ for step (i) and (iv), $(1/n)^{|Y|}(1-1/n)^{n-|Y|}$ for step (ii), at least $1/2$ for step (iii) by Lemma 11, and $(1/n)^{k-|Y|}(1-1/n)^{n-(k-|Y|)}$ for step (v), noting that here exactly the bits in $X\setminus Y$ have to be flipped. Multiplying these probabilities, we see that $\Pr[E_{Y}]\geq\tfrac{1}{2}(1-1/n)^{2n-k}\mu^{-2}n^{-k}\geq\tfrac{1}{2}e^{-2}\mu^{-2}n^{-k}$. There are $2^{k}$ disjoint events $E_{Y}$, each ensuring that after two iterations the population contains an individual $z$ as desired. Hence the probability $p$ to obtain such a $z$ within two iterations is at least $p\geq\sum_{Y}\Pr[E_{Y}]\geq 2^{k}\cdot\tfrac{1}{2}e^{-2}\mu^{-2}n^{-k}$.
By Lemma 12, the population always contains an individual satisfying the assumptions on $x$ and having the same objective value as $x$. Hence the time $T$ to generate the desired individual $z$ is stochastically dominated by two times a geometric distribution with success rate $p$, hence $E[T]\leq 2/p\leq 4e^{2}\mu^{2}n^{k}/2^{k}$. ∎
Since the proofs of Lemmas 6 and 7 for the classic SMS-EMOA mostly relied on elementary properties of standard bit-wise mutation and on the survival guarantee of Lemma 4, we can now use the survival guarantee of Lemma 12 to obtain analogous results for the SMS-EMOA with stochastic selection (at the price of requiring essentially twice the population size). This yields the following estimates for the time to obtain at least one inner Pareto optimum and the time to cover all inner Pareto front points starting from a population with at least one inner Pareto optimum.
@@
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2\overline{M}+1$ to optimize the $m$OJZJ problem. Then
  • after at most $e\mu(mk/2)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
  • after another at most $en\mu(n^{\prime}-2k+1)^{m/2}$ iterations in expectation, all inner Pareto front points are covered.
Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered. This analysis follows precisely the proof of Lemma 8 for the original SMS-EMOA except that we replace the estimate for expected time to obtain an extremal block, $e\mu n^{k}$ in that lemma, by the potentially lower number $e\mu n^{k}\min\{1,4e\mu/2^{k}\}$. This immediately gives the following result.
Lemma 16.
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2\overline{M}+1$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $e\mu((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})n^{k}\min\{1,4e\mu/2^{k}\}$ iterations in expectation, the full Pareto front is covered.
Combining Lemmas 15 and 16, we have the runtime of the SMS-EMOA with stochastic population update in the following theorem. It includes the case of two objectives, and then, with $\overline{M}=M=(n-2k+3)$, improves the previous result for this case, Theorem 2 in [BZLQ23], in particular by replacing the $2^{k/4}$ term by $2^{k}$.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq 2\overline{M}+1$ and stochastic population update to optimize the $m$OJZJ problem. Then after at most $e\mu(mk/2)^{k}(1+\ln m)+{en\mu(n^{\prime}-2k+1)^{m/2}+e\mu((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})n^{k}\min\{1,4e\mu/2^{k}\}}=O(\mu Mn^{k}\min\{1,4e\mu/2^{k}\})$ iterations in expectation, the full Pareto front is covered.
Comparing the runtime guarantees of Theorem 9 (classic SMS-EMOA) and Theorem 17 (SMS-EMOA with stochastic selection), we see that stochastic selection can at most lead to a speed-up by a factor of order $2^{k}/\mu$. Now $\mu$ is at least $\Omega(M)=\Omega((2n/m-2k+3)^{m/2})$. Consequently, the advantage of stochastic selection is rapidly decreasing with growing numbers of objectives and vanishes, e.g., when $m\geq k$.
6 Heavy-Tailed Mutation Helps
In the previous section, we saw that the advantage of the stochastic population update does not generalize well from bi-objective to many-objective optimization. We now regard another design choice that so far was only analyzed in bi-objective optimization, namely a heavy-tailed mutation operator. For this, we shall prove that the $k^{\Omega(k)}$ factor speed-up observed in bi-objective optimization extends to many objectives.
6.1 Heavy-Tailed Mutation
Different from the standard bit-wise mutation operator, which flips each bit independently with probability $1/n$, the heavy-tailed mutation operator proposed in [DLMN17] flips each bit independently with probability $\alpha/n$, where $\alpha$ follows a power-law distribution with parameter $\beta$. The number $\alpha$ is sampled anew in each application of the heavy-tailed mutation operator. The underlying power-law is defined as follows.
@@ 6.1 Heavy-Tailed Mutation
Let $n\in\mathbb{N}$ and $\beta>1$. We say that $\alpha$ follows a power-law distribution with (negative) exponent $\beta$ if for all $i\in[1..n/2]$, we have $\Pr[\alpha=i]=\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}i^{-\beta}$, where $C_{n/2}^{\beta}=\sum_{j=1}^{n/2}j^{-\beta}$.
The power-law in the choice of $\alpha$ extends to the offspring distribution in the way that an offspring has Hamming distance $j$ from the parent with probability $\Omega(j^{-\beta})$ [DLMN17]. This facilitates larger jumps in the search space, and thus escaping from a local optimum. We recall that for standard bit-wise mutation, the offspring has Hamming distance $j$ from the parent with probability is $j^{-\Omega(j)}$ only.
Since this observation was formulated in [DLMN17, Lemma 5.1] in a slightly impractical manner (with implicit constants independent of $\beta$, but only applying to distances in $(\beta-1,n/2]\cup\{1\}$), we first give a formulation that is more useful for most purposes. It treats $\beta$ as a constant, which makes a lot of sense in the light of all previous results, which gave the best performances for constant values of $\beta$.
Lemma 19.
Let $\beta>1$. Then there is a $c>0$ such that the following holds. Let $n\in\mathbb{N}$. Let $x\in\{0,1\}^{n}$ and $y$ be generated by applying the heavy-tailed mutation (with parameter $\beta$) to $x$. Let $H(x,y)$ be the Hamming distance between $x$ and $y$, let $j\in[1..n/2]$, and let $P_{j}^{\beta}:=P_{j}^{\beta}(n):=\Pr[H(x,y)=j]$. Then
$P_{j}^{\beta}\geq cj^{-{\beta}}.$
Proof.
If $\beta-1>\sqrt{n}$, then any $c>0$ such that $c\leq\min\{P_{j}^{\beta}(n)\mid n\in[1..(\beta-1)^{2}],j\in[1..n/2]\}$ trivially satisfied our claim. Hence, in the following, let $\beta-1\leq\sqrt{n}$.
If $j\in(\beta-1,n/2]\cup\{1\}$, then [DLMN17, Lemma 5.1 (1) and (3)] gives our result. If $j\in(1,\beta-1]$, then $j\leq\beta-1\leq\sqrt{n}$ and we conclude from [ZD23d, Lemma 16] that $P_{j}^{\beta}\geq\frac{\beta-1}{2\sqrt{2\pi}e^{8\sqrt{2}+13}\beta}j^{-\beta}$. Hence $c=\frac{\beta-1}{2\sqrt{2\pi}e^{8\sqrt{2}+13}\beta}$ suffices in this case. ∎
The heavy-tailed mutation operator has resulted in asymptotic performance gains by a factor of $k^{\Omega(k)}$ for the $(1+1)$ EA optimizing single-objective (classic and generalized) Jump functions with gap size $k$ [DLMN17, BBD24]. For the bi-objective OJZJ benchmark, again a speed-up of $k^{\Omega(k)}$ was proven when optimized via the GSEMO [ZD23c] and the NSGA-II [DQ23a]. Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choices [FQW18, WQT18, QGWF21, COY21, DELQ22, ABD22, DR23, DGI24].
@@ 6.2 Runtime
With proof ideas similar to those in Lemma 6 to 8, we prove the following three runtime estimates. The main differences are the different probabilities for generating a certain solution. However, these require somewhat different computations in the following lemma, which therefore appears slightly technical. We note that we did not optimize the result here, but were content with arriving at a bound that is small compared to the time needed to generate the extremal Pareto optima.
Lemma 21.
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with heavy-tailed mutation and with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem. Then after at most $O\mathopen{}\mathclose{{}\left((em/2)^{k}k^{\beta}\mu\ln m}\right)$ (for $m\geq 4$) or $O\mathopen{}\mathclose{{}\left(\binom{n}{k}k^{\beta}\mu/(n-k+1)}\right)$ (for $m=2$) iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
We use the same general as in Lemma 6 except for the following probability estimates. From Lemma 19, we know that the probability to generate a $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ from $x$ with $|x_{B_{i}}|_{1}\in[0..k-1]$ is at least
$\Omega\mathopen{}\mathclose{{}\left(\frac{\binom{n^{\prime}-|x_{B_{i}}|_{1}}{k-|x_{B_{i}}|_{1}}(k-|x_{B_{i}}|_{1})^{-\beta}}{\binom{n}{k-|x_{B_{i}}|_{1}}}}\right).$ (2)
Let first $m\geq 4$. For all $j\in[0..k-1]$, we compute
$\begin{split}&\mathopen{}\mathclose{{}\left(\frac{\binom{n^{\prime}-(j+1)}{k-(j+1)}}{\binom{n}{k-(j+1)}(k-(j+1))^{\beta}}}\right)\bigg{/}\mathopen{}\mathclose{{}\left(\frac{\binom{n^{\prime}-j}{k-j}}{\binom{n}{k-j}(k-j)^{\beta}}}\right)\\ ={}&{}\mathopen{}\mathclose{{}\left(\frac{(n^{\prime}-(j+1))!(k-(j+1))!(n-k+j+1)!}{(k-(j+1))!(n^{\prime}-k)!n!(k-(j+1))^{\beta}}}\right)\\ &{}\bigg{/}\mathopen{}\mathclose{{}\left(\frac{(n^{\prime}-j)!(k-j)!(n-k+j)!}{(k-j)!(n^{\prime}-k)!n!(k-j)^{\beta}}}\right)\\ ={}&{}\frac{(n^{\prime}-(j+1))!(n-k+j+1)!(k-j)^{\beta}}{(n^{\prime}-j)!(n-k+j)!(k-(j+1))^{\beta}}=\frac{n-k+j+1}{n^{\prime}-j}\mathopen{}\mathclose{{}\left(\frac{k-j}{k-(j+1)}}\right)^{\beta}\\ \geq{}&{}\frac{n-k+1}{n^{\prime}}\geq\frac{mn^{\prime}/2-n^{\prime}/2+1}{n^{\prime}}>1,\end{split}$ (3)
where the antepenultimate inequality uses $j\geq 0$, and the last inequality uses $m\geq 4$. Hence the expression in (2) is smallest for $|x_{B_{i}}|_{1}=0$, and hence at least of the order of
$\frac{\binom{n^{\prime}}{k}}{\binom{n}{k}}k^{-\beta}\geq\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{en}}\right)^{k}k^{-\beta}=\mathopen{}\mathclose{{}\left(\frac{2}{em}}\right)^{k}k^{-\beta}.$ (4)
For $m=2$, we have $n=n^{\prime}$. Since
$\binom{n-j}{k-j}=\binom{n-j}{n-k}\geq\binom{n-(k-1)}{n-k}=n-k+1$
for $j\in[0..k-1]$, and
$\frac{n-k+1}{\binom{n}{k-|x_{B_{i}}|_{1}}}(k-|x_{B_{i}}|_{1})^{-\beta}\geq\frac{n-k+1}{\binom{n}{k}}k^{-\beta},$ (5)
we can estimate the expression in (2) by $\Omega\mathopen{}\mathclose{{}\left((n-k+1)k^{-\beta}/\binom{n}{k}}\right)$.
Replacing (1) by (4) or (5) in the proof of Lemma 6, we obtain that the expected number of iterations to reach an inner Pareto optimum is $O\mathopen{}\mathclose{{}\left((em/2)^{k}k^{\beta}\mu\ln m}\right)$ for $m\geq 4$ and is $O\mathopen{}\mathclose{{}\left(\binom{n}{k}k^{\beta}\mu/(n-k+1)}\right)$ for $m=2$. ∎
Lemma 22.
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with heavy-tailed mutation and with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $\frac{e\beta}{\beta-1}n\mu(n^{\prime}-2k+1)^{m/2}$ iterations in expectation, all inner Pareto front points are covered.
Proof.
We use similar arguments as in the proof of Lemma 7 except for the following probability calculation. The probability to generate an offspring with function value equal to the one of an uncovered inner Pareto front point from a neighbor covered by the population is at least $\frac{1}{n\mu}P_{1}^{\beta}\geq\frac{\beta-1}{e\beta n\mu}$, where $P_{1}^{\beta}\geq\frac{\beta-1}{e\beta}$ is from [ZD23d, Lemma 16]. ∎
Lemma 23.
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with heavy-tailed mutation and with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $\binom{n}{k}C_{n/2}^{\beta}\mu k^{\beta}((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})$ iterations in expectation, the full Pareto front is covered.
Proof.
We reuse the arguments of Lemma 8 except for the following probability calculation. From Lemma 19, we know that the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\binom{n}{k}^{-1}P_{k}^{\beta}=\Omega\mathopen{}\mathclose{{}\left(\binom{n}{k}^{-1}\frac{1}{\mu}k^{-\beta}}\right).\qed$
Considering the whole process for the full coverage of the Pareto front from Lemmas 21 to 23, we have the following runtime result for the SMS-EMOA with heavy-tailed mutation (where we estimated $\binom{n}{k}\leq n^{k}/(k!)$ and $k!\geq\sqrt{2\pi}k^{k+0.5}e^{-k}$).
@@
Let $k\leq n^{\prime}/2$ and $\beta>1$. Consider using SMS-EMOA with $\mu\geq\overline{M}$ and with heavy-tailed mutation to optimize the $m$OJZJ problem. Then after at most $O(M\mu k^{\beta-0.5-k}(en)^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
@@
Compared to the runtime guarantee of Theorem 9 for the original SMS-EMOA, the guarantee of Theorem 24 above for the SMS-EMOA with heavy-tailed mutation is by a factor of asymptotically $k^{k+0.5-\beta}/e^{k}$ stronger. Compared to the SMS-EMOA with stochastic population update (Theorem 17), the speed-up is by a factor of $\min\{k^{k+0.5-\beta}/e^{k},4\mu k^{0.5-\beta}(k/2)^{k}/e^{k-1}\}$, which is super-exponential w.r.t. $k$.
7 Runtime Analysis for Two Classic Bi-objective Benchmarks